OpenStudy (anonymous):

Find the area of a regular hexagon with the given measurement. 2 sqrt 3 apothem A =

OpenStudy (anonymous):

@TheSmartOne @satellite73 @midhun.madhu1987

TheSmartOne (thesmartone):

\(\huge\color{green}{Area}=\color{orange}{\frac{1}{2}}\times\color{blue}{Perimeter}\times\color{red}{Apothem}\)

OpenStudy (anonymous):

thnx

TheSmartOne (thesmartone):

no prob.